# The Timescales of Embedded Star Formation as Observed in STARFORGE

### v1

@@ 4.3 Tracking the Evolution of Formation Time and Embeddedness for Individual Stars
We present these results in Figure 7, which shows the value of as a function of stellar age, in four simulation snapshots; a movie for all snapshots is available as supplementary material555We also include a link to the online movie here: https://zenodo.org/records/17156032. Each star is plotted as a point that is colored and scaled in size by the star’s maximum mass, using the same color bar as Figure 1, such that a star keeps a consistent color and size across all plots. The gray vertical line indicates zero age, when a star’s host sink particle reaches maximum mass and stops accreting; stars to the left of this line (negative age) are accreting sink particles. The grey area on the right of each panel indicates times longer than the run-time of the simulation at each timestep.
The histograms on the top and right axes track the bulk distributions in stellar age and , respectively. The top plots all stars (blue), and the subset of massive stars (; purple). The vertical gold line in the histogram (and associated triangle) indicates the median age of the stars that exist at that timestep, which is typically much less than the simulation age (c.f., the median age at is only 1.35 Myr). These histograms reinforce the general trend shown in Figure 4 — extended star formation and a shift towards lower extinction with increasing time.
We first focus on the overall timing of star formation throughout the simulation, best seen in the age histograms on the top of each plot. As discussed in previous STARFORGE works (e.g., Grudić et al., 2022; Millstone et al., 2023), stars form across essentially the entire duration of the simulation. There is a clear peak at , but it is both preceded and followed by on-going star formation, as can best be seen in the blue histogram for the oldest timestep (bottom right). This extended star formation is consistent with persistence of residual dense gas, as was clearly visible in the snapshot images in Figure 4.
We now turn to the joint distributions of age and extinction, as traced by individual stars. Throughout the simulation, young stars first appear in dense gas with high characteristic , consistent with forming inside of dense molecular gas. However, over time they become orders of magnitude less obscured, typically with a notable transition near the time they stop accreting mass, at the zero-age boundary. This pattern repeats for stars formed throughout the simulation. There is also a more gradual, but steady decrease in the typical extinction of the post-accretion () stars, due to the on-going loss of dense dusty gas throughout the cloud.
@@ 4.3.1 Mass-dependence of
4.3.2 What is Driving the Mass-dependence of ?
@@
The behavior of the low-mass stars is quite different, however. These stars will have limited ability to drive stellar feedback, given that the available energy during the pre-main sequence phase is strongly mass dependent (Section 4.2), and their final luminosities are low. It therefore is unlikely that their reduction in embeddedness is feedback driven.
However, low-mass stars are not entirely passive in shaping their local environments. Even though their radiative output is minimal, these stars can drive powerful collimated outflows during the protostellar phase, which have been shown to efficiently expel nearby gas on small scales (e.g., Offner & Chaban, 2017). However, the momentum input from low-mass protostars is substantially lower than that from high-mass stars, meaning that any reduction in extinction proceeds more slowly and may not reach large enough scales to fully de-embed the star. Indeed, in Figure 8, we see the evolution of extinction for low-mass stars takes place over a longer timescale, with no strong features in associated with reaching their maximum mass.
@@
4.4 How is the Time Evolution of Embdeddedness Impacted by Cloud Mass?
@@
Here, we investigate the empirical mass-dependence of the timescale for cluster embedding. As discussed in Section 4.1, the fiducial STARFORGE simulation is based on a M⊙ cloud, in radius. The free-fall time depends on the initial cloud density as , which for the fiducial model corresponds to . We now compare the timescales of the fiducial simulation to a second, more massive cloud simulation beginning with M⊙ and a radius. The initial density of this higher mass simulation is a factor of 2.7 smaller than the low mass simulation, leading to a 1.6 longer free-fall time for the massive cloud. Both simulations have identical initial virial parameters, such that their crossing times at the characteristic velocity dispersions scale linearly with the free-fall time. We therefore expect that the more massive cloud should take longer to collapse, and should have gas available for accretion over a longer timescale.
We test this assumption in Figure 9 by generating plots similar to the bottom of Figure 4 for both the low mass (red) and high mass (purple) simulations. The filled points indicate the time evolution of the fraction of stars with , which for both simulations starts at one (all stars highly obscured) and declines to small values (little net extinction) by the end of the simulation. The solid lines indicate the fraction of high mass () stars at each timestep (same accounting as right hand panel in Figure 6666We note that we could have chosen to measure the fraction of high mass stars which have reached their maximum mass, consistent with the age zero from Section 3.1. If we had made this distinction, in the Fiducial model increases slightly to 1.72 .) as a function of time, which increases from zero to one for both the low mass (red) and high mass (purple) cluster simulations, and decreases as stars explode and undergo supernovae.
@@
5 Observational Measures of Embeddedness
@@ 5.1 Infrared Emission
5.1.1 Total Dust Luminosity
@@
Using the same framework used to make the images, and calculate the dust luminosity in Section 5.1, we make mock 24 observations of the simulation using skirt. Here, we adopt the Spitzer 24 pass-band which ranges from 20.8 to 26.1 microns, integrating the dust SED only across these wavelengths to determine the 24 luminosity, as observed by Spitzer.
@@
Overall, the evolution in Figure 11 is quite similar to that seen for the overall dust luminosity. However, while the luminosity increases gradually as more stars form like the overall dust luminosity, the luminosity peaks earlier, at 6.2 Myr (1.68 ) compared to the peak at 6.8 Myr seen in Section 5.1. Additionally, in contrast to the relatively smooth decline in luminosity seen in in Figure 10, the decline in luminosity from the peak is much more pronounced. In fact, by 6.8 Myr when the total dust luminosity is peaking, the luminosity is already at 70% of its maximum.
@@ 5.2 Gas Density and H Emission
We note that, surprisingly, the strongest H emission actually precedes the formation of a significant ionized gas component in the cloud. As seen in Equation 3, recombination-driven H depends on the square of the gas density, which is highest earlier in the simulation. At later times, there is more ionized gas, but most of that gas is now diffuse, with a much lower recombination rate, and thus a lower H luminosity.
5.3 Optical and UV Emission from Stars
As gas and dust are cleared by stellar feedback, the direct emission from stars begins to escape, allowing optical and ultraviolet light to emerge. These bands not only mark the transition from embedded to exposed phases of stellar formation, but also offer critical age constraints on the system. To evaluate how this starlight evolves and eventually becomes observable, we begin by computing the bolometric luminosity of each star, based on its mass and radius at each evolutionary stage (Section 4.2).
@@ 5.3 Optical and UV Emission from Stars
To account for extinction, we determine the dust attenuation using the viewing angle-averaged values from Section 3.2. Extinction in each band is then scaled according to the Cardelli et al. (1989) extinction law, with UV extinction taken to be .
5.4 The Relative Timing of Observational Indicators
@@
To do this comparison, we consider quantities discussed throughout this section; dust luminosity (Section 5.1), H luminosity (Section 5.2), and the stellar luminosity as viewed in the optical and UV (Section 5.3. We present this comparison for potentially observed luminosities, where each quantity is measure through the specific telescope band pass with the corresponding wavelength attenuation.
@@
This unexpected behavior for the time-evolution of H and 24 emission again highlights the importance of local properties within the evolving cloud. Both luminosities will be strongly enhanced by having dense gas close to massive stars, although for different physical mechanisms – i.e., recombination versus higher dust temperatures, to a degree that frequently dominates over the global evolution of the majority of the gas777Note the degree to which the high surface brightness 24 emission in Figure 11 is actually highly localized in small clumps, likely tracing associated young bright stars..
@@
6 Discussion
6.1 Contextualizing STARFORGE and Observations
@@
6.1.1 The Timing and Sequence of Star Formation
The model of star formation presented in the STARFORGE simulations indicates that stellar associations are not necessarily single age populations. We measure stars forming throughout all times of the simulations, with a standard deviation of 2.14Myr and an rms of 6.7 Myr Section 3.1). Moreover, there is a clear mass progression in star formation, with low mass stars forming first, up to 1 Myr before high mass stars (8Myr) begin to form. We also note that some of the extended star formation is localized to a long-lasting and/or late-forming high-density gas region in the turbulent, but largely dissociated molecular gas cloud (Section 3.1.2).
However, most observational studies of embedded timescales typically assume stars form nearly instantaneously (e.g., Portegies Zwart et al., 2010; Hannon et al., 2022; Whitmore et al., 2023; Rodríguez et al., 2025). However, this assumption is increasingly challenged across a wide range of mass and spatial scales. In the most comprehensive studies of individual clusters like 30 Doradus (e.g., De Marchi et al., 2011), or Orion (e.g., Bally, 2008; Kounkel et al., 2018), often find evidence for extended star formation. Additionally, age spreads remain common when considering smaller associations and clusters (e.g., Kerr et al., 2024). Similarly, the extended star formation seen in STARFORGE echoes this broader observational trend, and is advocated in other theoretical works (e.g., Krumholz & McKee, 2020). Together, these results suggest that the assumption of near instantaneous star formation may be an oversimplification in many systems.
While the single age stellar population is not an appropriate model for the behavior of the STARFORGE simulations as a whole, it is true that some of the extended star formation is spatially segregated from the earliest star formation (Section 3.1.2, Guszejnov et al. (2022)). In an observational context, these sub-regions might be analyzed as independent regions, for which single age assumptions might be more appropriate approximations. However, in distant galaxies, these sorts of sub-regions are indistinguishable with current resolution, and cannot be analyzed independently. In such cases, it may be worth considering SED and spectral fits that allow a more extended or complex recent star formation history. Additionally, the simulations presented here use a limited set of initial conditions, and it may well be that the simultaneous, synchronous star formation assumption is more appropriate in other conditions.
@@ 6.1.1 The Timing and Sequence of Star Formation
Also pertaining to the sequence of star formation is the specific timing of the massive star formation, where quasi-instantaneous massive star formation is used in most timescale measurements (Schinnerer & Leroy, 2024). However, the timing of massive star formation is currently an open question in the field (e.g., Kumar et al., 2006; Rosen et al., 2020). The massive stars forming after the smaller stars observed here has also been seen in other simulations (Vázquez-Semadeni et al., 2024). This timing supports the mode of formation where smaller cores form first, and high mass cores are a scaled up version of these smaller cores.
6.1.2 Interpreting Measures of Embeddedness for Individual Stars
In analyzing the STARFORGE simulations, we identified two main pathways through which the embeddedness evolves. The first is the more traditional large scale gas clearance from the association (Section 4.1), driven by massive stars; within STARFORGE, this process has been analyzed in detail by Farias et al. (2024) across a range of cloud properties. The second is through highly localized feedback in the immediate neighborhood of individual massive stars (Section 4.3). This latter process acts such that as embedded massive stars reach their maximum mass and luminosity (i.e., ending the accretion phase of their formation), their typical extinction drops dramatically, within less than 1Myr. The emergence of low mass stars from the dust is much more gradual, more closely tracking the overall evolution of the cloud and their drift from regions of high gas density, rather than local feedback.
@@ 6.1.2 Interpreting Measures of Embeddedness for Individual Stars
6.1.3 Interpreting Measures of Embeddedness for Semi/Un-resolved Association and Star Forming Regions
In the absence of resolved stellar populations, clusters and associations are typically analyzed through their integrated spectral energy distributions. In STARFORGE, we find the expected sequence of emission that moves from luminous in the mid-IR, to a mixture of H and mid-IR emission, to a brief window where optical stellar luminosity is comparable to the other tracers before the optical and UV light becomes dominant (Section 5.4. The timescales associated with these transitions are largely in agreement with those measured in external galaxies (Chevance et al., 2022; Kim et al., 2021; Linden et al., 2024; Rodríguez et al., 2025; Knutas et al., 2025; Ramambason et al., 2025).

### v2

@@ IV.3 Tracking the Evolution of Formation Time and Embeddedness for Individual Stars
We present these results in Figure 7, which shows the value of as a function of stellar age, in four simulation snapshots; a movie for all snapshots is available as supplementary material666We also include a link to the online movie here: https://zenodo.org/records/17156032. Each star is plotted as a point that is colored and scaled in size by the star’s maximum mass, using the same color bar as Figure 1, such that a star keeps a consistent color and size across all plots. The dashed vertical line indicates zero age, when a star’s host sink particle reaches maximum mass and stops accreting; stars to the left of this line (negative age) are accreting sink particles. The grey area on the right of each panel indicates times longer than the run-time of the simulation at each timestep.
The density distribution histograms on the top and right axes track the bulk distributions in stellar age and , respectively. The top plots all stars (blue), and the subset of massive stars (; purple). The vertical gold line in the histogram (and associated triangle) indicates the median age of the stars that exist at that timestep, which is typically much less than the simulation age (c.f., the median age at is only 1.35 Myr). These histograms reinforce the general trend shown in Figure 4 — extended star formation and a shift towards lower extinction with increasing time.
We first focus on the overall timing of star formation throughout the simulation, best seen in the age histograms on the top of each plot. As discussed in previous STARFORGE works (e.g., Grudić et al., 2022; Millstone et al., 2023) and in Section III.1, stars form across essentially the entire duration of the simulation. There is a clear peak at , but it is both preceded and followed by on-going star formation, as can best be seen in the blue histogram for the oldest timestep (bottom right, as well as Figure 1). This extended star formation is consistent with persistence of residual dense gas, as was clearly visible in the snapshot images in Figure 4.
Top histograms are number counts for stars of different ages with blue being all stars and purple representing stars larger than 8 M⊙. Side histograms are number counts of stars at a given extinction. We see star formation occurring across the entire duration of the simulation, and high mass stars form throughout. The movie can also be found at https://doi.org/10.5281/zenodo.17156032.
We now turn to the joint distributions of age and extinction, as traced by individual stars. We encourage readers to view the movie version of Figure 7, where one can track each star in higher snapshot fidelity than is possible in the 4 panel still shots. Throughout the simulation, young stars first appear in dense gas with high characteristic , consistent with forming inside of dense molecular gas. Therefore, as stars form, they appear in the upper left corner of the plot. However, over time they become orders of magnitude less obscured (moving down the y-axis), typically with a notable transition near the time they stop accreting mass, at the zero-age boundary. This pattern repeats for stars formed throughout the simulation. There is also a more gradual, but steady decrease in the typical extinction of the post-accretion () stars, due to the on-going loss of dense dusty gas throughout the cloud.
@@ IV.3.1 Mass-dependence of
IV.3.2 What is Driving the Mass-dependence of ?
@@
The behavior of the low-mass stars is quite different, however. These stars will have limited ability to drive stellar feedback, given that the available energy during the pre-main sequence phase is strongly mass dependent (Section IV.2), and their final luminosities are low. It therefore is unlikely that their reduction in embeddedness is feedback driven.
However, low-mass stars are not entirely passive in shaping their local environments. Even though their radiative output is minimal, these stars can drive powerful collimated outflows during the protostellar phase (e.g., Frank et al., 2014), which have been shown to efficiently expel nearby gas on small scales (e.g., Offner & Chaban, 2017). Because the outflow momentum and radiative output are directly to the stellar mass (Offner & Chaban, 2017), low-mass protostars inject substantially less momentum per unit time than their high-mass counterparts, meaning that any reduction in extinction proceeds more slowly and may not reach large enough scales to fully de-embed the star. Indeed, in Figure 8, we see the evolution of extinction for low-mass stars takes place over a longer timescale, with no strong features in associated with reaching their maximum mass.
@@
IV.4 How is the Time Evolution of Embdeddedness Impacted by Cloud Mass?
@@
Here, we investigate the empirical mass-dependence of the timescale for cluster embedding. As discussed in Section IV.1, the fiducial STARFORGE simulation is based on a M⊙ cloud, in radius. The free-fall time depends on the initial cloud density as , which for the fiducial model corresponds to 3.7 Myr. We now compare the timescales of the fiducial simulation to a second, more massive cloud simulation beginning with M⊙ and a radius. The initial density of this higher mass simulation is a factor of 2.7 smaller than the low mass simulation, leading to a 1.6 longer free-fall time for the massive cloud. Both simulations have identical initial virial parameters, such that their crossing times at the characteristic velocity dispersions scale linearly with the free-fall time. We therefore expect that the more massive cloud should take longer to collapse, and should have gas available for accretion over a longer timescale.
We test this assumption in Figure 9 by generating plots similar to the bottom of Figure 4 for both the low mass (red) and high mass (purple) simulations. The filled points indicate the time evolution of the fraction of stars with , which for both simulations starts at one (all stars highly obscured) and declines to small values (little net extinction) by the end of the simulation. The solid lines indicate the fraction of high mass () stars at each timestep (same accounting as right hand panel in Figure 6777We note that we could have chosen to measure the fraction of high mass stars which have reached their maximum mass, consistent with the age zero from Section III.1. If we had made this distinction, in the Fiducial model increases slightly to 1.72 .) as a function of time, which increases from zero to one for both the low mass (red) and high mass (purple) cluster simulations, and decreases as stars explode and undergo supernovae.
@@
V Observational Measures of Embeddedness
@@
In the previous section, this framework led us to conclude that individual stars transition from deeply embedded to exposed very rapidly once they reach their maximum mass, and that this transition is driven primarily by local feedback from massive stars rather than by global gas removal (Section IV). To connect this physical picture to what observers actually measure, we must determine whether such locally driven emergence would be imprinted in the diagnostics used to infer embedded lifetimes. We therefore model the standard tracers used to infer embedded timescales and star–formation rates from our simulation, and measure how long each tracer is dominant, peaks, and fades relative to the physically defined de–embedding of the stars presented in the previous section. By comparing these synthetic timescales to those inferred from real systems, we can test both whether the short, local emergence phase around massive stars leaves a clear imprint on the luminosity evolution and whether the STARFORGE model is capturing the same physics that governs observed embedded clusters.
@@ V.1 Infrared Emission
V.1.1 Total Dust Luminosity
@@
Using the same framework used to make the images, and calculate the dust luminosity in Section V.1, we make mock 24 observations of the simulation using skirt. Here, we adopt the Spitzer 24 pass-band which ranges from 20.8 to 26.1 microns, integrating the dust SED only across these wavelengths to determine the 24 luminosity, as observed by Spitzer.
@@
Overall, the evolution in Figure 11 is quite similar to that seen for the overall dust luminosity. However, while the luminosity increases gradually as more stars form just like the overall dust luminosity, the luminosity peaks earlier, at 6.2 Myr (1.68 ) compared to the peak at 6.8 Myr seen in Section V.1. Additionally, in contrast to the relatively smooth decline in luminosity seen in in Figure 10, the decline in luminosity from the peak is much more pronounced. In fact, by 6.8 Myr when the total dust luminosity is peaking, the luminosity is already at 70% of its maximum.
@@ V.2 Gas Density and H Emission
We note that, surprisingly, the strongest H emission actually precedes the formation of a significant ionized gas component in the cloud. As seen in Equation 3, recombination-driven H depends on the square of the gas density, which is highest earlier in the simulation. At later times, there is more ionized gas, but most of that gas is now diffuse, with a much lower recombination rate, and thus a lower H luminosity. This behavior echoes the picture developed in Section IV.2, where H is brightest when massive stars are still tightly coupled to localized gas, and fades once their feedback evacuates their immediate surroundings, even as global ionization of the cloud continues to grow.
V.3 Optical and UV Emission from Stars
As gas and dust are cleared by stellar feedback, the direct emission from stars begins to escape, allowing optical and ultraviolet light to emerge. These bands not only mark the transition from embedded to exposed phases of stellar formation, but also offer critical age constraints on the system. To evaluate how this starlight evolves and eventually becomes observable, we begin by computing the bolometric luminosity of each star, based on its mass and radius at each evolutionary stage (Section IV.2).
@@ V.3 Optical and UV Emission from Stars
To account for extinction, we determine the dust attenuation using the viewing angle-averaged values from Section III.2. Extinction in each band is then scaled according to the Cardelli et al. (1989) extinction law, with UV extinction taken to be .
V.4 The Relative Timing of Observational Indicators
@@
To do this comparison, we consider quantities discussed throughout this section; dust luminosity (Section V.1), H luminosity (Section V.2), and the stellar luminosity as viewed in the optical and UV (Section V.3. We present this comparison for potentially observed luminosities, where each quantity is measure through the specific telescope band pass with the corresponding wavelength attenuation.
@@
This unexpected behavior for the time-evolution of H and 24 emission again highlights the importance of local properties within the evolving cloud. Both luminosities will be strongly enhanced by having dense gas close to massive stars, although for different physical mechanisms – i.e., recombination versus higher dust temperatures, to a degree that frequently dominates over the global evolution of the majority of the gas888Note the degree to which the high surface brightness 24 emission in Figure 11 is actually highly localized in small clumps, likely tracing associated young bright stars..
@@
VI Discussion
VI.1 Contextualizing STARFORGE and Observations
@@
VI.1.1 The Timing and Sequence of Star Formation
The model of star formation presented in the STARFORGE simulations indicates that stellar associations are not necessarily single age populations. We measure stars forming throughout all times of the simulations, with a standard deviation of 2.14Myr and an rms of 6.7 Myr Section III.1). Moreover, there is a clear mass progression in star formation, with low mass stars forming first, up to 1 Myr before high mass stars (8Myr) begin to form. We also note that some of the extended star formation is localized to a long-lasting and/or late-forming high-density gas region in the turbulent, but largely dissociated molecular gas cloud (Section III.1.2).
However, most observational studies of embedded timescales typically assume stars form nearly instantaneously (e.g., Portegies Zwart et al., 2010; Hannon et al., 2022; Whitmore et al., 2023; Rodríguez et al., 2025). However, this assumption is increasingly challenged across a wide range of mass and spatial scales. In the most comprehensive studies of individual clusters like 30 Doradus (e.g., De Marchi et al., 2011), or Orion (e.g., Bally, 2008; Kounkel et al., 2018), often find evidence for extended star formation. Additionally, age spreads remain common when considering smaller associations and clusters (e.g., Kerr et al., 2024). Similarly, the extended star formation seen in STARFORGE echoes this broader observational trend, and is advocated in other theoretical works in both turbulence supported clouds (e.g., Krumholz & McKee, 2020), and collapse driven models (e.g., Vázquez-Semadeni et al., 2024b). Together, these results suggest that the assumption of near instantaneous star formation may be an oversimplification in many systems.
While the single age stellar population is not an appropriate model for the behavior of the STARFORGE simulations as a whole, it is true that some of the extended star formation is spatially segregated from the earliest star formation (Section III.1.2, Guszejnov et al. (2022)). In an observational context, these sub-regions might be analyzed as independent regions, for which single age assumptions might be more appropriate approximations. However, in distant galaxies, these sorts of sub-regions are indistinguishable with current resolution, and cannot be analyzed independently. In such cases, it may be worth considering SED and spectral fits that allow a more extended or complex recent star formation history. Additionally, the simulations presented here use a limited set of initial conditions, and it may well be that the simultaneous, synchronous star formation assumption is more appropriate in other conditions.
@@ VI.1.1 The Timing and Sequence of Star Formation
Also pertaining to the sequence of star formation is the specific timing of the massive star formation, where quasi-instantaneous massive star formation is used in most timescale measurements (Schinnerer & Leroy, 2024). However, the timing of massive star formation is currently an open question in the field (e.g., Kumar et al., 2006; Rosen et al., 2020). The massive stars forming after the smaller stars observed here has also been seen in other simulations (Vázquez-Semadeni et al., 2024a). This timing supports the mode of formation where smaller cores form first, and high mass cores are a scaled up version of these smaller cores.
VI.1.2 Interpreting Measures of Embeddedness for Individual Stars
In analyzing the STARFORGE simulations, we identified two main pathways through which the embeddedness evolves. The first is the more traditional large scale gas clearance from the association (Section IV.1), driven by massive stars; within STARFORGE, this process has been analyzed in detail by Farias et al. (2024) across a range of cloud properties. The second is through highly localized feedback in the immediate neighborhood of individual massive stars (Section IV.3). This latter process acts such that as embedded massive stars reach their maximum mass and luminosity (i.e., ending the accretion phase of their formation), their typical extinction drops dramatically, within less than 1Myr. The emergence of low mass stars from the dust is much more gradual, more closely tracking the overall evolution of the cloud and their drift from regions of high gas density, rather than local feedback.
@@ VI.1.2 Interpreting Measures of Embeddedness for Individual Stars
VI.1.3 Interpreting Measures of Embeddedness for Semi/Un-resolved Association and Star Forming Regions
In the absence of resolved stellar populations, clusters and associations are typically analyzed through their integrated spectral energy distributions. In STARFORGE, we find the expected sequence of emission that moves from luminous in the mid-IR, to a mixture of H and mid-IR emission, to a brief window where optical stellar luminosity is comparable to the other tracers before the optical and UV light becomes dominant (Section V.4. The timescales associated with these transitions are largely in agreement with those measured in external galaxies (Kim et al., 2021; Linden et al., 2024; Rodríguez et al., 2025; Knutas et al., 2025; Ramambason et al., 2025).
This agreement suggests that the same physical processes that govern light emergence in our simulations here are the same as those in the observed clusters. In particular, the simulations reveal that de-embedding is a highly local process, driven primarily by the feedback from massive stars disrupting their immediate surroundings, rather than the result of global cloud dispersal. Because such fine-grained spatial information is inaccessible in unresolved systems, the results presented here serve as a valuable calibration tool, offering physical insight into what integrated tracers are actually measuring. Observed timescales for light emergence can therefore be interpreted as reflecting the timing and locality of massive star formation, rather than a synchronized, global clearance of dust and gas.